OpenStudy (anonymous):

Solve the equation. 408 = s - 129 A. s = 537 B. s = 622 C. s = 735 D. s = 800

OpenStudy (anonymous):

A

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

im ina fwist gwade

OpenStudy (anonymous):

he first thing you want to do is get s on one side and everything else on the other side. 408 = s - 129 (To do this, you have to add 129 to both sides) 408 + 129 = s - 129 + 129 537 = s

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

oops typo....the first thing ....(lol)

OpenStudy (anonymous):

It is A......100% positive.